40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 1 00:05:03 2013, Koji, Update, LSC, More stable lock of PRMI (REF33I and AS55Q) Screenshot.png
    Reply  Wed May 1 10:42:35 2013, Steve, Update, LSC, cleanup  
    Reply  Wed May 1 14:42:45 2013, Jenne, Update, LSC, POP now has lens in front of PD 
Message ID: 8517     Entry time: Wed May 1 00:05:03 2013     Reply to this: 8518   8519
Author: Koji 
Type: Update 
Category: LSC 
Subject: More stable lock of PRMI (REF33I and AS55Q) 

[Jenne Koji]

- Today the spots were moving more than the usual. The OPLEV screens showed that the spots are too much off from the center.

- Each vertex OPLEVs were checked and OPLEV wonderland was discovered: Other than the usual misalignment of the spots,
it was found that PRM/ITMX/ITMY beams were clipped somewhere in the paths, BS/PRM oplevs had many loose components
including the input lenses (they are still clamped by a single dog clamp THIS SHOULD BE FIXED ASAP).

- On the ITMY table there were so many stray optics. They were removed and put on the wagon next to the ITMY table.
THIS SHOULD BE CLEARED ON THE WEDNESDAY CLEANING SESSION.

- During this OPLEV session, LSCoffset nulling was run.

- After the OPLEV session, the locking became really instantaneous. We wonder which of the OPLEV cleaning, LSC offset nulling,
and the usual seismic activity decay in the evening was effective to make it better.

- Initially the lock was attempted with REFL33I/Q and some ~10sec lock streches were obtained. During this lock,
  the optical gain of AS55Q was measured in relative to REFL33Q. In deed they were calibrated to be the same
  gain at the input matrix.

- After the MICH signal source was switched to AS55Q, the lock streches became more regular and the minutes long.
We precisely tuned the phase of AS55 and REFL55 in terms of the differential excitation of ITMX/Y using lockin (FREQ 250, AMP 1000).

- We noticed that the AS port spot with AS55Q MICH was darker than the REFL33Q MICH. This suggests the existence of residual offset
in REFL33Q. In deed we observed +30cnt offset in REFL33Q when the PRMI is locked with AS55Q MICH.

- Phases and relative gains of the signals were as follows:

PRCL: REFL33I 1.00 =REFL55I +0.4
MICH: AS55Q 29deg x1.00 = REFL33Q -14deg x1.00 = REFL55Q 118deg 0.03?

- We tried to lock PRMI with AS55Q. The acquisition was not as easy as that with REFL33I. This might be from the saturation of the
REFL55I signal. This configuration should tested with different whitening gain. Handing off using the input matrix went well once the
lock was obtained by REFL33I.

- Handing off from AS55Q to REFL55Q was not successful.

- At the end of the session, Jenne told me that the POP PD still has a large diameter beam. (and a steering mirror with a peculiar reflection angle.)
==> THIS SHOULD BE FIXED ASAP
because the normalization factor can be too much susceptible to the misalignment of the spot.

- The configuration of the filters:

PRCL FM3/4/5/6 G=+0.05 / NORM 0.04 POP110I
MICH FM4/5 G=-5.00 / NORM 0.01 POP110I (or none)

Screenshot.png

ELOG V3.1.3-